Gant Accounting performs two types of services, Audit and Tax. Gant’s overhead costs consist of computer support, $312000; and legal support, $156000. Information on the two services is:

  Audit   Tax  
Direct labor cost $50000 $100000
CPU minutes 40000 10000
Legal hours used 200 800


Overhead applied to audit services using activity-based costing is

$187200.

$156000.

$312000.

$280800.
Answer #1

Ans.

There are two overheads namely computer support and legal support. Computer support will be allocated between audit amd tax in ratio of CPU minutes and legal support will be allocated in ratio of legal hours.

Calculation of overhead applied to audit services

Computer support= 312000*40000/50000

= 249600

Legal support=156000*200/1000

=31200

Total overhead= 249600+31200

= $280800